OpenStudy (anonymous):

PLEASE HELP on jan 1, 2010, Chessville has a population of 50,000 people. Chessville then enters a period of population growth. Its population increases 7% each year. On the same day, Checkersville has a population of 70,000 people. Checkersville starts to experience a population decline. its population decrease 4% each year. During what year will the population Chessville first exceed that of Checkersville? Show work and explain steps.

OpenStudy (stamp):

What have you worked so far?

OpenStudy (anonymous):

50,000 * .07x= ?

OpenStudy (anonymous):

and 70000 * .04x= ?

OpenStudy (stamp):

Do you have a textbook? Look up expontential growth and decay, or half life in your index and find the page.

OpenStudy (anonymous):

no i dont have a text book with me. staying at my dads. forgot it. all i remember about is that you take them and sorta make em into 1 equation

OpenStudy (anonymous):

50,000(1.07)^t = 70,000(1-.04)^t = 70,000(.96)^t (1.07/.96)^t= (70,000/50,000)=1.4 t ln (1.07/.96) = ln(1.4) .10848 t = .33647 t=3.102 So population will be equal in 3.102 years

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

ur welcome

OpenStudy (darrius):

WOah when did mik get smart lolol jk

OpenStudy (anonymous):

shhhh im not smart but ik this one lmfao

OpenStudy (anonymous):

Do I write it exactly like u did or just put the answer?

OpenStudy (anonymous):

I have this same question

OpenStudy (anonymous):

@mikala1 can you explain how you did it?